# A quick probability-oriented introduction to operator splitting methods

### v1

@@ 4. The Trotter-Kato formula
Multiplicative methods that feature a fixed one-step internal optimization step (such as the Marchuk-Yanenko splitting) can be called alternatively locally one-dimensional methods (LOD) in a generalized sense [188], while the Peaceman-Rachford and Douglas-Rachford splittings are known as alternating direction implicit methods (ADI).
@@
the numbers $\{a_{k}\},\{b_{k,i,j}\}$ not necessarily being non-negative or even real. Such methods are often used to construct higher order schemes. However, the term ”exponential splitting” sometimes means just the Lie-Trotter splitting.
@@
([188]) Using (5.7), one can define
@@
The above is a modern classification of operator splitting methods presented in [148]. However, multiplicative and additive methods historically have different origins and different societies of contributors; a specific operator splitting scheme or a particular version of a general scheme (e.g. the Störmet-Verlet integration method) may have been developed or rediscovered as a narrow ad hoc method for a particular problem, so the nomenclature is rather diverse and does not necessarily possess integrity a researcher interested in historical accuracy would desire262626even though not to the degree of W. Rudin’s ironic remark [293]: ”Thus it appears that Čech proved the Tychonoff theorem, whereas Tychonoff found the Čech compactification – a good illustration of the historical reliability of mathematical nomenclature.”.
@@
We start with the Lie-Trotter splitting. A standard and historically justified alternative name is the fractional step method (or the method of fractional steps), which can also include other similar schemes. Other possibilities revolve around (sur)names of S. Lie, T.E. Trotter and T. Kato in different combinations, one of most common being the Lie splitting [148]. Analysts often speak about the Trotter-Kato formula only. Probabilists may use ”splitting-up”. Physicists may speak about the Suzuki-Trotter expansion/decomposition. Sequential splitting is another alternative.
@@
The autonomous Peaceman-Rachford and Marchuk-Yanenko splittings can be written in the terms of revolvents. Thus the sequential splitting of (5) and the resolvent splitting of e.g. (5.2), (5.3) or (5.5) can be considered separately. Another example is the Ryu splitting (e.g. [13, 244]).
@@
Particular versions and algorithms include the Störmet-Verlet and leapfrog methods (see Section 10), time-evolving block decimation [326, 179], split Hamiltonian272727in particular, split Hamiltonian MC methods [70, 302], split-step Fourier [8, 321], mapping method [337, 243], gradient-projection [92, 130], split Bregman methods [152, 272], primal-dual splitting [48, 81], Rosenbrock’s approximate matrix factorization [49, 29], the aforementioned Tseng, Davis-Lin, Ryu splittings and others.
@@
Discretization here may involve basis expansions, interpolation between points of a space(-time) grid, choosing a specific formula for the first/second/higher derivative at a boundary, variable grid steps, finite differences and finite volumes in general, using volume and mass preservation formulations etc. Alternatively, if a probabilistic interpretation is available, one may apply MC methods, and the upper bound of the error becomes, roughly speaking,
@@
where $E_{MC}$ is a statistical discrepancy (due to an insufficient number of simulations), estimated e.g. with using the Berry-Esseen, Bikelis or concentration inequalities [162], and $E_{sim}$ is the error of the actual simulation procedure for the corresponding random process.
@@
(6.4) ${S}^{\varepsilon}_{h}\Big{(}\frac{1}{N}\sum_{k=\overline{1,N}}u_{0}(\xi_{x,n,k})\Big{)},$
respectively. The error of these numerical schemes comes from a number of sources:
@@
On the other hand, real life engineering PDEs are usually stated for bounded domains automatically.
@@
Though MC is significantly less sensitive to the problem of unbounded domains, the presence of a boundary can cause problems. One can think about using the classical Euler-Maruyama scheme to simulate a one-dimensional SDE whose solution always stays positive: it is impossible to keep the approximations positive since they use symmetrical random increments, so one has to consider so called tamed schemes etc. to remedy this.
@@
The boundary $\partial D$ is assumed to be constant and piecewise Lipschitz, as is customary.
We can apply the Feynman-Kac formula to get the analogs of (6). $a$ is now defined only on $D,$ so one could try to consider a decomposition
@@
and a family of degenerate first order PDEs to define $\widetilde{S}_{t}\circ g=:\widetilde{S}_{t}^{g}$
@@
instead of (6.2).
Then (6) transform into
@@
(6.6) $\left(\widetilde{S}_{h}\circ T_{h}\right)u_{0}(x)$ $=\widetilde{S}_{h}\Big{(}\operatorname{E}u_{0}(\xi_{x}(h))\!\!\,{\textbf{1}}\left[\tau>h\right]+\operatorname{E}\widehat{u}_{b}\left(\xi_{x}(\tau)\right)\!\!\,{\textbf{1}}\left[\tau\leq h\right]\Big{)},$
@@
Even though we consider only the Dirichlet BC, it is clear that the decomposition (6.5) cannot be arbitrary even if $f$ can accommodate it. E.g. if $\sigma=0$ on $C\subset\partial D,$ these parts $C$ may be unreachable for $\xi$ (so BC does not make sense here at all) or the process $\xi$ cannot be reflected there (so no Neumann BCs are allowed) etc. Recalling that no classification of boundary points of a diffusion exists in dimensions higher than or equal to $2,$ we can see this as a particular difficulty of the probabilistic origin.
@@
where the term $g_{\varepsilon}$ appears due to boundary corrections when calculating derivatives and explicitly depends on $\widehat{u}_{b},$ and $v_{\varepsilon,0}$ is a discretized version of an IC (which changes at each step of the splitting scheme). Let ${v}_{\varepsilon,n}$ denote finite-dimensional approximations of $(T_{t})_{t\geq 0}$ on the grid $G_{\varepsilon}$ at time step $n.$ One can use the standard $\theta-$scheme [^29] This is the Crank-Nicolson scheme if $\theta=\frac{1}{2}.$ with time step $h$ for solving ODEs to get
@@
which should be solved for ${v}_{\varepsilon,n+1}$. As it has been remarked already, the matrix $B_{\varepsilon}$ depends on $G_{\varepsilon}$ and contains the second power of the grid size of $G_{\varepsilon}.$ Thus it needs implicit and sufficiently stable methods whose step sizes get smaller as $\varepsilon\to 0$ and thus limit the choice of a grid in practice. This introduces an error the value of which depends on $G_{\varepsilon},$ the domain $D,$ properties of $\partial D$ etc. Such schemes are run at every step of the splitting procedure.
@@
$\frac{1}{N}\sum_{k=\overline{1,N}}J^{\varepsilon}\circ\widetilde{S}^{\varepsilon}_{h}\circ I^{\varepsilon}\circ u^{\varepsilon}_{0}(\xi_{x,n,k}),$
(6.10) $J^{\varepsilon}\circ\widetilde{S}^{\varepsilon}_{h}\circ I^{\varepsilon}\Big{(}\frac{1}{N}\sum_{k=\overline{1,N}}u^{\varepsilon}_{0}(\xi_{x,n,k})\Big{)},$
@@
    the decomposition of a BC and the corresponding corrections (if any);
@@
    internal time discretization.
We will continue the discussion of how BCs affects splitting later.
@@
All these schemes – (6), (6), (6), (6), (6) – are different interpretations of the same splitting procedure. Regardless of the path, the end result shows the discrepancy between a theoretical analysis of an abstract splitting scheme and a practical implementation in this particular case, and the same discrepancy is present for any abstract splitting method if the implementation process is ignored. If a Faedo-Galerkin scheme is used, (2) is replaced with properties of the scheme, bases and spaces involved etc.
@@
Actually, one can initially use approximate discretized semigroups to formulate a splitting scheme. This is how matrix-valued splitting schemes appear. [280] can be consulted for such a formulation in the case of standard approximation schemes and corresponding applications of the Trotter-Kato theorems and e.g. [200, 27, 26, 119, 217] can be consulted for abstract results for splitting methods in this setting.
Again, a particular splitting scheme can be selected first in practice, and optimization subroutines may follow naturally. Then errors of different parts of the pipeline can be treated separately.
@@
In the case of a general PDE and a general method (e.g. the finite element method), an additional matrix may appear in (6) before the time derivative [188], so even the ”explicit” forward Euler method of (6) would use matrix inversion, which requires some additional effort.
@@
A separation of linear and nonlinear parts for a Cauchy problem (e.g. for advection-diffusion-reaction equations) is often a viable strategy (see e.g. [188, 299, 148, 98, 96] and references therein) in general due to technical and numerical considerations and may even lead to theoretical results by directly exploiting properties of the flow of the ODE that corresponds to the nonlinear part (e.g. [123] and Example 21). For instance, this ODE may be solvable explicitly.
@@
An important and frequent task is to find a steady-state/equilibrium solution with $\frac{\partial u}{\partial t}=0.$ Such solutions often represent the long-term behavior of a physical system. In general, operator splitting methods act on a fixed time interval, and it is advised to use other methods to seek such solutions or to provide a modified version of a splitting scheme [148, 47, 188].
@@ 7. Some remarks on the relation between ACPs and practice
These two intertwined frameworks compliment each other. For instance, a variational reformulation of the Trotter-Kato theorem can be found in [200]. One can recall two different formulations of SPDEs as an example of the interplay between these two competing frameworks [238, 141].
@@
Obviously, there are a difference between theoretical difficulties and engineering challenges [^30] For instance, such an extreme pathology as the Lebesgue thorn of the potential theory cannot appear in real life. and a gap between available rigorous theoretical results and practical needs. However, applications is exactly the role operator splitting methods excel in besides theoretical studies. In fact, the development of additive operator splitting methods and semi-discretized multiplicative ones was motivated and driven by practical studies.
@@
It is nontrivial to find the accuracy of a splitting scheme in a particular case. We have already discussed the approach that combines the local error and the stability estimates. The reader can consult Section 5 for many references which typically use the theory of semigroups. To obtain the rate of convergence, nontrivial additional assumptions should be added.
Some results on the speed of convergence (and occasionally on directly related topics such as stability preservation) that have not been mentioned yet earlier and are at least to some degree outside the analytic setting of the classical Kato-Trotter formula are [114, 113, 124, 125, 192, 193, 305, 117, 118, 276, 178, 276, 220, 174, 319, 20, 192, 196, 195, 104, 105, 306, 140, 145, 186, 275, 210, 146, 225, 180, 315, 316, 317, 318, 315, 127, 42, 39, 23, 27, 26, 115, 212, 172, 87]. See also references later in this section.
@@
In particular: [192, 193] study the difference between the Feynman-Kac formula and the Shrödinger operator using probabilistic techniques including the theory of Levý processes (see also references therein); [319, 20, 192, 196] obtain estimates on the difference of the corresponding kernels, also via probabilistic techniques.
As illustrations, we include the following results about applications (and occasionally about implementation) of splitting methods: [^31] We do not include pure simulations. We try to give a wide range of applications or methods and sometimes prefer recent publications since they can easy lead to older results in case a reader is interested. By an application, a specific equation may be understood. [112, 299, 203, 52, 307, 281, 145, 144, 143, 208, 184, 147, 207, 69, 9, 10, 24, 156, 183, 249, 173, 235, 79, 236, 111, 237, 17, 348, 213, 217].
In general, the rate of the convergence can be arbitrary (or a bound may not be available at all). E.g. [236] establishes only the well-posedness of a scheme and its energy-preservation properties (cf. [261]).
@@
[235, 111, 237, 17] provide a bound of the speed of the convergence that combines sizes of time and space discretizations.
@@
([342]) Let $A_{1},A_{2},A=A_{1}+A_{2}$ be non-negative self-adjoint operators. If
@@
This covers the example of two Laplacians on a bounded domain with a smooth boundary, one with a Dirichlet BC and the other with a Neumann BC, and $\alpha\in(\frac{1}{2};\tfrac{3}{4}).$
@@
([198, 197], cf. [344]) Let $A_{1},A_{2}$ be non-negative self-adjoint operators such that $A=A_{1}+A_{2}$ is self-adjoint on $D(A_{1})\cap D(A_{2}).$ Then
@@
([174], see also [276, 124, 117, 118]) Consider an ACP on a Hilbert space
@@
See also the survey [88].
Though most references above and from the previous sections are heavily analytic, it is not obligatory to use only such methods and a variety of only vaguely related to the semigroup theory323232which is itself a wide field or totally different approaches are used.
@@
([119], cf. [323]; also see [235, 111, 237, 17, 172, 217]) A shortened exposition of a general variational framework for one semi-discrete multiplicative splitting scheme with the backward Euler internal step is as follows. Assume that separable reflexive Banach spaces $V_{j},j=\overline{0,m},\cap_{j=\overline{1,m}}=V_{0},$ are continuously and densely embedded into a Hilbert space $H$ so one can consider Gelfand triples
@@
Additionally we assume that the norms $\|\|_{V_{0}}$ and $\sum_{j=\overline{1,m}}\|\|_{V_{j}}$ are equivalent. For fixed $j,$ let $A_{j}(t),t\in[0;T],$ be hemicontinuous coercive monotone333333see Section 5 operators from $V_{j}$ to $V_{j}^{*}$ such that $\sum_{j=\overline{0,m}}A_{j}(t)=A_{0}(t),t\in[0;T],$ and for some $p>1$
@@
$v_{n,k}\approx u_{0}\left(\frac{k}{n}\right),\quad k=\overline{0,n}.$
@@
The general analysis of parabolic problems in domains with arbitrary BCs seems to be missing and assumptions of general theorems should be checked in every particular case and are not guaranteed to apply. Note that a BC goes directly into the definition of the corresponding domain.
Now we present a short illustrative description of one very different method of studying errors of a splitting scheme which has not been represented in the previous references (with a few exceptions).
@@
([188], cf. e.g. [37, 127, 305]) In the autonomous one-dimensional case this idea gives for the Lie-Trotter splitting with time step $h$
@@
([341]) Consider the Strang splitting $S_{2}(\tfrac{h}{2})S_{1}(h)S_{2}(\tfrac{h}{2})$ for
@@
A related and often applied simultaneously approach is based on establishing a PDE for the global or local errors. Such PDEs, being treated as IVPs or ODEs in functional spaces, usually lead to expressions for the errors in the terms of the variation-of-constants formula and again involve multiple iterated commutators of differential operators. Alternatively, an error can be given as a difference of integral representations of the exact solution and the approximation. This formally works for almost any operators but one still needs to prove the validity of such formulae and obtain bounds for them.
@@
([306]) The error of the Lie-Trotter splitting is

### v2

@@ 4. The Trotter-Kato formula
Multiplicative methods that feature a fixed one-step internal optimization step (such as the Marchuk-Yanenko splitting) can be called alternatively locally one-dimensional methods (LOD) in a generalized sense [HundsVer07Numerical], while the Peaceman-Rachford and Douglas-Rachford splittings are known as alternating direction implicit methods (ADI).
@@
the numbers $\{a_{k}\},\{b_{k,i,j}\}$ not necessarily being non-negative or even real. Such methods are often used to construct higher order schemes. However, the term ”exponential splitting” sometimes simply means the standard Lie-Trotter splitting.
@@
([HundsVer07Numerical]) Using (5.7), one can define
@@
The above is a modern classification of operator splitting methods presented in [GlowOshYin17Splitting]. However, multiplicative and additive methods historically have different origins and different societies of contributors; a specific operator splitting scheme or a particular version of a general scheme (e.g. the Störmet-Verlet integration method) may have been developed or rediscovered as a narrow ad hoc method for a particular problem, so the nomenclature is rather diverse and does not necessarily possess integrity a researcher interested in historical accuracy would desire282828even though not to the degree of W. Rudin’s ironic remark [Ru91functional]: ”Thus it appears that Čech proved the Tychonoff theorem, whereas Tychonoff found the Čech compactification – a good illustration of the historical reliability of mathematical nomenclature.”.
@@
We start with the Lie-Trotter splitting. A standard and historically justified alternative name is the fractional step method (or the method of fractional steps), which can also include other similar schemes. Other possibilities revolve around (sur)names of S. Lie, T.E. Trotter and T. Kato in different combinations, one of most common being the Lie splitting [GlowOshYin17Splitting]. Analysts often speak about the Trotter-Kato formula only. Probabilists may use ”splitting-up”. Physicists may speak about the Suzuki-Trotter expansion/decomposition. Sequential splitting is another alternative.
@@
The autonomous Peaceman-Rachford and Marchuk-Yanenko splittings can be written in the terms of revolvents. Thus the sequential splitting of (5) and the resolvent splitting of e.g. (5.2), (5.3) or (5.5) can be considered separately. Another example is the Ryu splitting (e.g. [ArCamTam21strengthened, MaTam23resolvent]).
@@
Particular versions and algorithms include the Störmet-Verlet and leapfrog methods (see Section 10), time-evolving block decimation [UrSol16parallel, HaHaMcCu20hybrid], split Hamiltonian292929in particular, split Hamiltonian MC methods [CaSanzShaw22split, ShanLanJohnNeal14split], split-step Fourier [AleZaShei05split-step, TaSmithWolf02analysis], mapping method [Wis82origin, Mal94mapping], gradient-projection [CuiShanb16analysis, FiNoWright07gradient], split Bregman methods [GolBreOsher10geometric, OberOsherTaTsai11numerical], primal-dual splitting [BotCsetHend14recent, ConKiContHi23proximal], Rosenbrock’s approximate matrix factorization [Botver03new, BeckGonPeWei14comparison], the aforementioned Tseng, Davis-Lin, Ryu splittings and others.
@@
Discretization here may involve choosing a mesh, basis expansions, interpolation between points of a space(-time) grid, choosing a specific formula for the first/second/higher derivative at a boundary, variable grid steps, finite differences and finite volumes in general, using volume and mass preservation formulations etc. Alternatively, if a probabilistic interpretation is available, one may apply MC methods, and the upper bound of the error becomes, roughly speaking,
@@
where $E_{MC}$ is a statistical discrepancy (due to an insufficient number of simulations), estimated e.g. with using the Berry-Esseen, Bikelis or concentration inequalities [GraTa13stochastic], and $E_{sim}$ is the error of the actual simulation procedure for the corresponding random process.
@@
(6.4) ${S}^{\varepsilon}_{h}\Big{(}\frac{1}{N}\sum_{k=\overline{1,N}}u_{0}(\xi_{\cdot,n,k})\Big{)}(x),$
respectively. The error in these numerical schemes comes from a number of sources:
@@
On the other hand, ”real-world” engineering PDEs are usually stated for bounded domains automatically.
@@
Though MC is significantly less sensitive to the problem of unbounded domains, the presence of a boundary can cause problems. One can think about using the classical Euler-Maruyama scheme to simulate a one-dimensional SDE whose solution always stays positive: it is impossible to keep the approximations positive since they use symmetrical random increments, so one has to consider corrected or suitable implicit schemes etc. to remedy this [^31] See e.g. [Alfon05discretization, DeNeuSzpruch12Euler] for results and references..
@@
The boundary $\partial D$ is assumed to be piecewise Lipschitz, as is customary.
We can apply the Feynman-Kac formula to get the analogs of (6). $f$ is now defined only on $D\times{\mathbb{R}},$ so one could try to consider a decomposition (cf. Section 9)
@@
and a family of degenerate first order PDEs [^32] For instance, one can add small noise to define this PDE properly. to define $\widetilde{S}_{t}\circ g=:\widetilde{S}_{t}^{g}$
@@
instead of (6.2). Since the BC is a function of $x$ only, an alternative is setting $\widehat{u}_{b}=\widetilde{u}_{b}=u_{b}.$
Remark 33.
However, references in Section 9 show that such naive approaches to BC are not proper ones.
Then (6) transforms into
@@
(6.6) $\left(\widetilde{S}_{h}\circ T_{h}\right)u_{0}(x)$ $=\widetilde{S}_{h}\circ\Big{(}\operatorname{E}u_{0}(\xi_{\cdot}(h))\!\!\,{\textbf{1}}\left[\tau>h\right]+\operatorname{E}\widehat{u}_{b}\left(\xi_{\cdot}(\tau)\right)\!\!\,{\textbf{1}}\left[\tau\leq h\right]\Big{)}(x),$
@@
Even though we consider only the Dirichlet BC, it is clear that the decomposition (6.5) cannot be arbitrary even if $f$ can accommodate it. E.g. if $\sigma=0$ on some $C\subset\partial D,$ this set $C$ may be unreachable for $\xi$ (so BC does not make sense here at all) or the process $\xi$ cannot be reflected there (so the Neumann BC cannot be imposed) etc. Recalling that no classification of boundary points of a diffusion exists in dimensions higher than or equal to $2,$ we can see this as a particular difficulty of the probabilistic origin.
@@
where the term $g_{\varepsilon}$ appears due to boundary corrections when calculating derivatives and explicitly depends on $\widehat{u}_{b},$ and $v_{\varepsilon,0}$ is a discretized version of the IC (which changes at each step of the splitting scheme). Let ${v}_{\varepsilon,n}$ denote finite-dimensional approximations of $(T_{t})_{t\geq 0}$ on the grid $G_{\varepsilon}$ at time step $n.$ One can use the standard $\theta-$scheme [^33] This is the Crank-Nicolson scheme if $\theta=\frac{1}{2}.$ with time step $h$ for solving ODEs to get
@@
which should be solved for ${v}_{\varepsilon,n+1}$. As it has been remarked already, the matrix $B_{\varepsilon}$ depends on $G_{\varepsilon}$ and contains the second power of the grid size of $G_{\varepsilon}.$ Thus it needs implicit and sufficiently stable methods whose step sizes get smaller as $\varepsilon\to 0$ and thus limit the choice of the grid in practice. This introduces an error the value of which depends on $G_{\varepsilon},$ the domain $D,$ properties of $\partial D$ etc. Such schemes are run at every step of the splitting procedure.
@@
$\frac{1}{N}\sum_{k=\overline{1,N}}J^{\varepsilon}\circ\widetilde{S}^{\varepsilon}_{h}\circ I^{\varepsilon}\circ u^{\varepsilon}_{0}(\xi_{\cdot,n,k})(x),$
(6.10) $J^{\varepsilon}\circ\widetilde{S}^{\varepsilon}_{h}\circ I^{\varepsilon}\Big{(}\frac{1}{N}\sum_{k=\overline{1,N}}u^{\varepsilon}_{0}(\xi_{\cdot,n,k})\Big{)}(x),$
@@
    the decomposition of BC and the corresponding corrections (if any);
@@
    internal time discretization (and associated internal errors).
We will continue the discussion of how BC affects splitting later.
@@
All these schemes – (6), (6), (6), (6), (6) – are different interpretations of the same splitting procedure. Regardless of the path, the end result shows the discrepancy between a theoretical analysis of our abstract splitting scheme and a practical implementation of it, and the same discrepancy is present between any abstract splitting method and its implementation. For instance, if a Faedo-Galerkin scheme is used, (2) is replaced with properties of the scheme, bases and spaces involved etc.
@@
Actually, one can initially use approximate discretized semigroups to formulate a splitting scheme. This is how matrix-valued splitting schemes appear. [Pazy12Semigroups] can be consulted for such a formulation in the case of standard approximation schemes and corresponding applications of the Trotter-Kato theorems and e.g. [ItoKa02evolution, BatCsoNi09operator_splittings, BaCsoFarNi12operator_spatial, EisenHan22variational, KoLu17numerical] can be consulted for abstract results for splitting methods in this setting.
Again, a particular splitting scheme can be selected first in practice, and optimization subroutines may follow naturally. Then different parts of the pipeline can be tuned separately.
@@
In the case of a general PDE and a general method (e.g. the finite element method), an additional matrix may appear in (6) before the time derivative [HundsVer07Numerical], so even the ”explicit” forward Euler method of (6) would use matrix inversion, which requires some additional effort.
@@
A separation of linear and nonlinear parts for a Cauchy problem (e.g. for advection-diffusion-reaction equations) is often a viable strategy (see e.g. [HundsVer07Numerical, Schatz02numerical, GlowOshYin17Splitting, DesDuDuLouMa11adaptive, DesDuDuLouMa14analysis] and references therein) in general due to technical and numerical considerations and may even lead to theoretical results by directly exploiting properties of the flow of the ODE that corresponds to the nonlinear part (e.g. [Faou09Analysis] and Example 21). For instance, this ODE may be solvable explicitly.
@@
An important and frequent task is to find a steady-state/equilibrium solution with $\frac{\partial u}{\partial t}=0.$ Such solutions often represent the long-term behavior of a physical system. In general, operator splitting methods act on a fixed time interval, and it is advised to use other methods to seek such solutions or to provide a modified version of a splitting scheme [GlowOshYin17Splitting, Bo23analysis, HundsVer07Numerical].
@@ 7. Some remarks on the relation between ACPs and practice
These two intertwined frameworks compliment each other. For instance, a variational reformulation of the Trotter-Kato theorem can be found in [ItoKa02evolution]. One can recall two different formulations of SPDEs as an example of the interplay between these two competing frameworks [LiuRock15Stochastic, GaMand11Stochastic].
@@
To conclude, we have a difference between theoretical difficulties and engineering challenges [^34] Such an extreme pathology as the Lebesgue thorn of the potential theory cannot appear in real life; theoretical technical advances on irregular and regular boundary points (in the sense of Wiener) for domains with corners are rarely of interest for pure engineers etc. and a gap between available rigorous theoretical results and practical needs. However, applications is exactly the role operator splitting methods excel in besides theoretical studies. In fact, the development of additive operator splitting methods and semi-discretized multiplicative ones was motivated and driven by practical studies.
@@
It is nontrivial to find the accuracy of a splitting scheme. We have already discussed the approach that combines the local error and the stability estimates. The reader can consult Sections 4-5 for many references which typically use the theory of semigroups. To obtain the rate of convergence, nontrivial additional assumptions should be added.
Some results on the speed of convergence (and occasionally on directly related topics such as stability preservation) that have not been mentioned yet earlier and are at least to some degree outside the analytic (semigroup-based) setting of the classical Kato-Trotter formula are [EiSchna18error, EiJahnSchnau19error, FaouOsSchratz15analysis, FaGei07iterative, IchiTa97estimate, IchiTa00estimate_2, SheAza20splitting, EinOs15overcoming_1, EinOs16overcoming_2, OsSchratz13stability, HanOs16high, OsSchratz13stability, KuhWa01commutator, HanOs08dimension, Ta97error, AzuIchi08note, IchiTa97estimate, IchiTa06exponential, IchiTa98error, DiaSchat96commutateurs, DiaSchatz9estimations, Sheng94Global, GaLaMon19unbalanced, Gei08fourth, HolKarlLie00operator, OsSchratz12error, KirWads02solution, Gei15recent, La16convergence, HaSu05finding, Su91general, Su92general, Su95hybrid, Su93improved, Su91general, FaHa07consistency, BlaCaMu08splitting, BlaCaCharMu13optimized, BatCsoEn12Stability, BatCsoNi09operator_splittings, BaCsoFarNi12operator_spatial, EinMoOs18efficient, KochLu08variational, HanHe17additive, CsoEhrFas21operator]. See also references later in this section. References given in Section 1 ([GlowOshYin17Splitting, GloLeTa89Augmented, Glo03Finite, Ya71Method, RyuYin23large_scale, BauCom17convex, Mar82methods, Mar88splitting, Mar86splitting] and others) can be used as a starting point for anyone interested in splitting methods in numerical methods (especially in the field of fluid mechanics and other physical applications and/or variational methods), and the corresponding literature is extremely rich.
@@
In particular: [IchiTa97estimate, IchiTa00estimate_2] study the difference between the Feynman-Kac formula and the Shrödinger operator using probabilistic techniques including the theory of Levý processes (see also references therein); [Ta97error, AzuIchi08note, IchiTa97estimate, IchiTa06exponential] obtain estimates on the difference of the corresponding kernels, also via probabilistic techniques.
As illustrations, we include the following results about applications (and occasionally about implementation) of splitting methods: [^35] We do not include pure simulations. We try to give a wide range of applications or methods and sometimes prefer recent publications since they can easy lead to older results in case a reader is interested. A specific important equation may be understood as an ”application”. [DuMaDes11parareal, Schatz02numerical, JahnAl10efficient, BouCar14splitting, ShiLiZhaoZou14new, PengZhaoFeng17operator, Gei08fourth, Gei11computing, Gei11iterative_book, KarlLieNatNordDah01operator, HolKarlLie00operator_2, GiRe19operator, Ka05accurate, CaiFangChenSun23fast, AltZi23secondArx, AltKoZi23bulk, BaCsoFar13operator, GonHerPe14Rosenbrok, HeLawDraPet14local, MarTeichWol16parabolic, HanKraOs12second, LiuMasRi23convergence, ChoiChoiKoh23convergence, LiuWangWang22second_order, DuanChenLiuWang22second_order, LiuWangWangWise22convergence, AuKaKochThal17convergence, Ze22analysis, KochLu11variational, KoLu17numerical].
In general, the rate of the convergence can be arbitrary (or a bound may not be available at all). E.g. [LiuWangWang22second_order] establishes only the well-posedness of a scheme and its energy-preservation properties (cf. [NeidZa99Trotter]).
@@
[LiuMasRi23convergence, DuanChenLiuWang22second_order, LiuWangWangWise22convergence, AuKaKochThal17convergence] provide a bound of the speed of the convergence that combines sizes of time and space discretizations.
@@
([Za05Trotter]) Let $A_{1},A_{2},A=A_{1}+A_{2}$ be non-negative self-adjoint operators. If
@@
This covers the example of two Laplacians on a bounded domain with a smooth boundary, one with a Dirichlet BC and the other with a Neumann BC ($\alpha\in(\frac{1}{2};\tfrac{3}{4})$).
@@
([IchiTaTaZa01note, IchiTa01norm], cf. [Zag22notes]) Let $A_{1},A_{2}$ be non-negative self-adjoint operators such that $A=A_{1}+A_{2}$ is self-adjoint on $D(A_{1})\cap D(A_{2}).$ Then
@@
([HanOs08dimension], see also [OsSchratz13stability, FaouOsSchratz15analysis, EinOs15overcoming_1, EinOs16overcoming_2]) Consider an ACP on a Hilbert space
@@
See also the survey [CsoSi21numerical].
Though most references above and from the previous sections are heavily analytic, it is not obligatory to follow this route and a variety of methods only vaguely related to the semigroup theory363636which is itself a wide field or totally different approaches are used.
@@
([EisenHan22variational], cf. [Te68stabilite]; also see [LiuMasRi23convergence, DuanChenLiuWang22second_order, LiuWangWangWise22convergence, AuKaKochThal17convergence, HanHe17additive, KoLu17numerical]) A shortened exposition of a general variational framework for one semi-discrete multiplicative splitting scheme with the backward Euler internal step is as follows. Assume that separable reflexive Banach spaces $V_{j},j=\overline{0,m},$ with $\cap_{j=\overline{1,m}}V_{j}=V_{0},$ are continuously and densely embedded into a Hilbert space $H$ so one can consider Gelfand triples
@@
The norms $\|\|_{V_{0}}$ and $\sum_{j=\overline{1,m}}\|\|_{V_{j}}$ are assumed to be equivalent. For fixed $j,$ let $A_{j}(t),t\in[0;T],$ be hemicontinuous coercive monotone373737see Section 5 operators from $V_{j}$ to $V_{j}^{*}$ such that $\sum_{j=\overline{0,m}}A_{j}=A_{0}$ on $[0;T],$ and for some $p>1$
@@
$v_{n,k}\approx u_{0}\Big{(}\frac{k}{n}\Big{)},\quad k=\overline{0,n}.$
@@
The general analysis of parabolic problems in domains with arbitrary BC is in development and assumptions of general theorems should be checked in every particular case and are not guaranteed to apply.
Now we present a short illustrative description of one very different method/tool of studying errors of a splitting scheme which has not been represented in the previous references (with a few exceptions).
@@
([HundsVer07Numerical], cf. e.g. [Bjor98operator, FaHa07consistency, SheAza20splitting]) In the autonomous one-dimensional case this idea gives for the Lie-Trotter splitting with time step $h$
@@
([Yo90construction]) Consider the Strang splitting $S_{2}(\tfrac{h}{2})S_{1}(h)S_{2}(\tfrac{h}{2})$ for
@@
A related and often used simultaneously approach is to derive PDEs for global or local errors. Such PDEs, being treated as IVPs or ODEs in functional spaces, usually lead to expressions for the errors in the terms of the variation-of-constants formula and again involve multiple iterated commutators of differential operators. Alternatively, an error can be given as a difference of integral representations of the exact solution and the approximation. This formally works for almost any operators but one still needs to prove the validity of such formulae and obtain bounds for them.
@@
([Sheng94Global]) The error of the Lie-Trotter splitting is